Calculate the volume of 0.02 mol L-1 potassium permanganate, KMnO4, solution that contains 1.0 X 10-3 mol of potassium permanganate.
1
Expert's answer
2020-08-02T15:52:39-0400

concentration = amount in solute/volume

0.02 = 1.0 *10-3/ volume

therefore volume = 1.0* 10-3/0.02

0.003/0.02 = 0.15 dm3